1. Consider the production function ?(?, ?) = (?1/2 + ?1/2)2/3 , where L denotes labor and K capital. This production function exhibits

A. constant returns to scale.

C. decreasing returns to scale.

D. increasing returns to scale.
Answer #1

Answer

C. decreasing returns to scale.

To find it we need to increase the input by n time sna check the output increases by how much time

0*qu = ( +7)*U=( +7)(u) = ((Yu)+(TU)) = mau Y + 7) = 0

It means the input increases by n times then the output increases by n^(1/3) times means the output increases less than an input so it is decreasing return to scale.